OpenStudy (anonymous):

How do the graphs of y = one over x and y = three over x – 4 compare?

OpenStudy (freckles):

y=1/x where does f(x)=1/x have a vertical asymptote?

OpenStudy (freckles):

What number can you not divide by?

OpenStudy (alexandervonhumboldt2):

y=1/x y=3/x -4